OpenStudy (anonymous):

Y=-x^2-10x-28 Vertex form

OpenStudy (mertsj):

Can you do it now that Satellite showed you how?

OpenStudy (anonymous):

I plugged your question into mathway.com. This is the answer it gave to the equation: y=^2-10x-28 Vertex (5,-53) Focus 5, 211/4 Direxrix y=-213/4 Axis of Symmetry x=5